# Semantic Edge Computing and Semantic Communications in 6G Networks: A Unifying Survey and Research Challenges

### v2

@@ 2.3 Unified Overview
Figure 3 overviews SEC and SemCom. SemCom consists of a DNN on the mobile device (transmitter) for semantic and channel coding and a DNN on the edge device (receiver) for decoding. The encoder and decoder DNNs are trained in an end-to-end fashion to maximize a semantic similarity metric between the source and destination messages. Meanwhile, the mobile DNN and edge DNN in SEC– which are used to extract the necessary information from the input and decode the information for inference – are equivalent to the encoder-decoder framework in SemCom. Indeed, TL-SemCom could be considered as a type of SEC that aims to achieve task-oriented communication by leveraging distributed DNNs.
As a result, many techniques in SEC to compress the semantic information without losing the task-oriented performance can be seamlessly adopted in SemCom to improve the communication efficiency. Similarly, one major advancement of SemCom is its capability to learn the optimal transmission scheme that is robust to wireless distortions by considering the corruption of the intermediate representations during training while SEC only utilizes conventional wireless technologies to transfer the semantics. Integrating these techniques from SemCom into SEC can achieve a lower end-to-end latency through a more robust data transmission link.
Distinction from existing surveys. Despite high relevance between these two fields, there is few work which focuses on the co-design of SEC and SemCom. Therefore, our goal is to provide a comprehensive review for both domains and hope this work can facilitate future investigations bridging these two fields. This sets our work apart from existing surveys such as [2, 33] that solely focus on the SemCom and [3] which exclusively covers SEC.
@@ 3.1 Taxonomy of Semantic Edge Computing
As discussed in Section 2.1, SEC primarily focuses on two problems: (1) how to optimize the computation of the mobile-edge system based on their resource constraints; and (2) how to extract meaningful semantics while compressing task-irrelevant information to reduce the transmission overhead. Based on these two directions, we categorize recent progress in SEC into system optimization and semantic compression. In system optimization, literature can be further categorize as static resource-aware optimization and dynamic model partitioning. In each category, DNNs can be distributed based on its sequential topology or Directed Acyclic Graph (DAG) topology. Based on the compression techniques investigated in literature, semantic compression can be further classified as dimension reduction, bit rate reduction and learning-driven rate-distortion optimization. Dimension reduction aims to introduce bottleneck layers at the splitting point of DNNs to compress the dimension of semantics, which can be further classified as reconstruction-based and task-oriented compression. Bit rate reduction leverages either lossless or lossy coding techniques to reduce the data size that need to be transmitted. The taxonomy of SEC is depicted in Figure 4.
@@ 3.1.1 System Optimization
The system optimization problem such as Eqn. 1 is often NP-hard, hence requiring advanced approaches to find the optimal solution. Based on the optimization approach, literature can be categorized as heuristic search [38], solver-based [34], relaxation-based [37] as well as Deep Reinforcement Learning (DRL)-based approaches [39]. Based on the partitioning graph, it can also be classified as sequential-based [40] and DAG-based [41] partitioning.
@@ 3.1.2 Semantic Compression
The transmission overhead in SEC can be reduced by compressing the dimension and bit-width of semantic representations. For example, a general approach to add an auto-encoder to encode the original intermediate representation to a lower space on mobile devices and reconstruct the representation at the edge. Also, knowledge distillation [42], Neural Architecture Search (NAS) [43] and neural network pruning [44] to manually create a bottleneck at the splitting point in order to reduce the dimensionality of latent representations. On the other hand, literature in SEC leverages different source coding approaches such as lossless coding (e.g. run-length coding [45], entropy coding [16]) and lossy coding (e.g. JPEG [46], quantization [6]) to further minimize the the average bit-length of latent representations. Recently inspired by the neural image compression [47], learning-driven rate-distortion optimization approaches are adopted to SEC [16]. The key idea is to optimize DNN based on the rate-distortion metric during training. After training, the learned latent representation can achieve minimum rate with entropy coding under certain distortion level. In addition, dynamic feature compression as an emerging technique has been studied recently to achieve wireless channel adaptive compression of latent representations [48].
@@ 3.2.1 Resource-Aware Partitioning
Sequential Partitioning. Eshratifar et al. [37] investigated layer-wise splitting to minimize the energy consumption of mobile devices and end-to-end latency. It first profiled the energy consumption and execution latency of each DNN layer during run-time. The optimal DNN partitioning is then converted into integer linear programming problem and lagrange relaxation based method [51] is used to constitute a lookup table for a different set of energy and networking constraints. Similarly, Zawish et al. [36] optimized the partitioning under various computing resource constraints (number of FLOPs), energy constraints and bandwidth constraints. Gao et al. [34] benchmarked layer-wise processing time on mobile and edge devices of different hyper parameters such as convolutional kernel size and data size. After benchmarking, a processing time estimation approach based on polynomial regression is developed. A mixed integer linear programming problem is formulated to minimize the end-to-end latency and solved with an opensource solver. Gao et al. [40] further considered the DNN partitioning in a multi-user scenario where multiple mobile devices will connect to the same edge device. The study jointly optimize the partitioning point and task scheduling to minimize inference latency, energy consumption and service fee paid to the server based on the aggregative game theory. Ding et al. [52] considered a privacy-preserving scenario for resource-limited DNN partitioning. The proposed approach first collects per-layer computation and memory cost to build a set of valid splitting points. Then a distance correlation metric [53] is leveraged during training to reduce the information leakage of latent representations. Li and Bi [39] extended the DNN partitioning to the multi-user cooperative inference system. The splitting and resource allocation are jointly optimized by transforming the problem into a mixed integer non-linear programming. A DRL algorithm is used to select splitting point to minimize energy consumption while a convex optimization algorithm is used for resource allocation.
@@
DAG Partitioning. While above-mentioned literature considered sequential splitting, Li et al. [41] investigated the DAG-based splitting and inference parallelism to accelerate the multiple inference requests. It involves partitioning the DNN model and offloading parts to cloudlets while using multi-threading to meet inference delay requirements. Duan and Wu [54] categorized DNN graphs as tree-structure and path-independent structure and proposed different algorithms regarding the type of DAG to optimize end-to-end latency under resource constraints. Similarly, Wang et al. [38] partitioned DNN based on its DAG structure to minimize latency. The problem is modeled with mixed integer programming which is proven np-hard and can be solved with heuristic search. Duan and Wu [55] extended the partitioning problem for multiple DNN tasks and proposed to minimize overall latency by jointly optimizing splitting point and task scheduling.
To optimize the computation and latency under resource constraints, most of the literature relies on an exhaustive performance analysis for every layer in DNNs [37, 34, 40, 52]. This raises a scalability challenge when DNN models become larger. Furthermore, different types of hardware that have distinct computing characteristics can result in different latency for the same computation overhead [56]. Only profiling layer-wise execution of DNNs on a single device may fail to generalize to heterogeneous platforms.
@@ 3.2.2 Dynamic Model Partitioning
Sequential Partitioning. Neurosurgeon [35] is one of the earliest work considering dynamic model partitioning in different bandwidth and energy constraints. It first leverages a regression model to estimate the computing load (number of FLOPs) of different DNN layers. Then, a greedy search approach is leveraged to select the splitting point in run-time to minimize the end-to-end latency or energy consumption of mobile devices. Li et al. [57] combined DNN splitting with early exits and optimized the splitting point and exit point jointly in the changing networking environment where the bandwidth varies drastically based on the bayesian online change point detection [58]. Mehta and Shorey [59] proposed to split DNN dynamically based on the changing computation resources to minimize the overall transmission consumption. However, [59] does not consider the dynamic networking condition. Karjee et al. [60] optimize the splitting point to minimize the computation latency in the dynamic computing resource constraints. In order to further optimize the communication latency, a dynamic network switching approach by leveraging both DRL and rule-based selection are proposed to change the wireless techniques such as bluetooth, cellular and Wi-Fi based on the network traffic. Lim and Kim [61] proposed to minimize the end-to-end latency using Lyapunov optimization. By modeling the dynamics in different time steps as a Lyapunov drift, [61] utilized an real-time splitting approach to jointly determine the DNN splitting point and the dynamic voltage and frequency scaling of the GPU on devices. Bakhtiarnia et al. [18] first profiled natural bottlenecks in off-the-shelf DNNs where the output is smaller than the original input data. Then, the DNN is dynamically allocated on mobile and edge device aiming at minimizing the end-to-end latency. Ko et al. [62] investigated the dynamic model partitioning in the distributed severless edge scenario where different head and tail models are running in containers with different computing status. The optimization problem is modeled with a Markov decision process and linear programming is used to choose the optimal container to minimize the inference latency. Nagamatsu and Hara-Azumi [63] utilized mixed-precision NAS [64] to optimize the bit rate of latent representations for different workloads. As the bit rate can be dynamically changed, the DNN is dynamically partitioned based on greedy search in order to minimize the inference time. Lee et al. [65] proposed a two-stage wireless channel adaptive model partitioning addressing the energy and memory constraint of devices. The splitting point is first determined using an off-line exhaustive search without considering the wireless channel. After deployment, the model can be adjusted online to minimize inference latency according to wireless channel conditions. This approach reduced the search space in the online adjustment phase making it easy to deploy in real time. However, the limited adjustment points can result in sub-optimal latency minimization.
@@
DAG Partitioning. Along with chain-based splitting, literature also investigate DNN partitioning based on their DAG structure. For example, Hu et al. [66] investigated the dynamic DAG partitioning in different DNN workload scenarios to minimize the latency in low network traffic scenario and maximize the overall throughput in heavy network traffic scenario. Zhang et al. [67] introduced a dynamic DAG partitioning algorithm that can adjust the splitting point to its nearly vertex in the run time. Liang et al. [68] further extended the work in [66] to a computing resource constrained scenario. Li et al. [69] proposed an adaptive DAG-based splitting to optimize the latency and resource utilization in the multi-user scenario.
The challenge of dynamic model partitioning is the cost of reconfiguration. In order to synchronize the deployment on both mobile and edge devices, additional orchestration is needed which result in extra communication and computation overhead. In addition, most of the dynamic partitioning algorithms rely on a greedy search [35, 63, 59] which can lead to large latency when the number of splitting points increases. Therefore, how to scale up the dynamic splitting to larger DNNs while maintaining a real-time performance needs to be further investigated.
@@
3.3 Advancements Semantic Compression
@@ 3.3.1 Dimension Reduction
On the other hand, inspired by DNN compression, the task-oriented bottleneck proposes to directly squeeze the network architecture and train the modified model with task-oriented loss (e.g. cross-entropy loss). Model compression approaches such as knowledge distillation [42], NAS [43] and DNN pruning [44] are often leveraged to achieve superior compression performance while preserving accuracy. For example, Eshratifar et al. [4] modified the size of the early layers in the DNN to decrease computational load on mobile devices. The bottleneck is introduced in the middle of the modified architecture and the model is trained with cross entropy loss. However, it suffers performance loss as introducing bottleneck in early layers can result in loss of task-oriented information. As such, Matsubara and Levorato [77] proposed to train the bottleneck using knowledge distillation. Similarly, Dong et al. [78] optimized the feature compression design using NAS while Guo et al. [79] utilized pruning to reduce the bottleneck size.
A common challenge in dimension reduction approaches is that large compression rate in early layers can jeopardize the final task performance. To overcome this, literature usually applies various training approaches to preserve the accuracy. For example, Yao et al. [73] applied multi-stage training to auto-encoder. In the first stage, the auto-encoder is trained to reconstruct the semantic features and in the second stage, knowledge distillation is applied to fine-tune the auto-encoder. Hao et al. [74] add a performance regularization term to reconstruction loss when training the auto-encoder. After training the auto-encoder, the partitioned DNN on the edge device is fine-tuned to further recover the performance. Matsubara et al. [15] proposed using knowledge distillation to train the light-weight DNN on mobile device (i.e., head network). Matsubara et al. [6] combined head knowledge distillation with fine-tuning to further mitigate the performance drop.
@@
[92] pruning information bottleneck Yes

### v3

@@ 2.4 Unified Overview
Although SEC and SemCom are derived independently, they share significant similarities in achieving fast and reliable edge intelligence. Figure 3 overviews SEC and SemCom. SemCom consists of a DNN on the mobile device (transmitter) for semantic and channel coding and a DNN on the edge device (receiver) for decoding. The encoder and decoder DNNs are trained in an end-to-end fashion to maximize a semantic similarity metric between the source and destination messages. Meanwhile, the mobile DNN and edge DNN in SEC– which are used to extract the necessary information from the input and decode the information for inference – are equivalent to the encoder-decoder framework in SemCom. Indeed, TL-SemCom could be considered as a type of SEC that aims to achieve task-oriented communication by leveraging distributed DNNs.
However, existing research in each field often overlooks the development in the other domain, resulting in suboptimal approaches for achieving intelligence in 6G networks. SemCom research typically disregards resource constraints in mobile devices and use a large encoder structure to perform more effective semantic extraction. For instance, [48] utilized a large ResNet backbone for semantic feature extraction on mobile devices while deploying only two linear layers for decoding at the edge server for tiny ImageNet classification tasks. While these approaches improve the inter-device communication efficiency, they significantly increase the computation overhead on the mobile device and do not make fully use of the computation power at the edge server. On the other hand, SEC methods such as [49, 50] only leverage existing wireless communication technologies to transmit latent representations. These methods require full-stack processing to achieve resilience in wireless channels which can introduce additional communication latency.
Therefore, we provide a comprehensive review of both domains and propose to incorporate SEC with SemCom across three dimensions: system level, application level, and physical level. From a system optimization perspective, the collaborative operation of DNNs across multiple devices involves heterogeneous computation capacity and network conditions. Therefore, this dimension focuses on optimizing computational and communication load on each device to achieve optimal inference latency and energy consumption. Many optimization methods in SEC such as linear programming [51] and Deep Reinforcement Learning (DRL) [52] can be applied to optimize the utilization of resources. In addition, system level optimization also encompasses the task scheduling and resource sharing to optimize the overall throughput which is missing in multi-view [53] or multi-user [54] SemCom.
Application level optimization encompasses the concept of semantic extraction and compression in SEC and SemCom, with its emphasis on the performance of DNN applications. The objective is to train DNNs that can effectively compress the semantic information to be sent while preserving task-oriented performance. Model compression methods, including knowledge distillation [55], Neural Architecture Search (NAS) [56] and pruning [57] can be applied to compress the data size. In addition, application level optimization can be enhanced through carefully designed loss functions that balance compression efficiency with task performance [58] or through multi-stage advanced training methodologies [6].
Physical level optimization involves a cross-layer design to overcome the physical layer distortions. Compared to current research in SEC that leverages existing wireless protocols, SemCom methods that directly encode semantic information for wireless transmission can substantially decrease end-to-end latency by bringing DNNs to the physical layer [59]. Such an approach requires less processing time to execute the full-stack protocol and enhance the reliability of communication by directly incorporating distortions during DNN training. In addition, physical level optimization involves adaptive transmission techniques [60, 61] that selectively transmit semantic information based on content importance and channel conditions can substantially enhance energy efficiency by reducing the transmission of non-critical data over unreliable channels.
Note that these three dimensions should not be optimized independently but require joint optimization to achieve optimal performance, representing a significant challenge in current state-of-the-art approaches. We hope our discussion can facilitate future investigations in integrating these design methodologies.
@@ 3.1 Taxonomy of Semantic Edge Computing
As discussed in Section 2.2, SEC primarily focuses on two problems: (1) how to optimize the computation of the mobile-edge system based on their resource constraints; and (2) how to extract meaningful semantics while compressing task-irrelevant information to reduce the transmission overhead. Based on these two directions, we categorize recent progress in SEC into system optimization and semantic compression. In system optimization, literature can be further categorized as static resource-aware optimization and dynamic model partitioning. In each category, DNNs can be distributed based on their sequential topology or Directed Acyclic Graph (DAG) topology. Based on the compression techniques investigated in the literature, semantic compression can be further classified as dimension reduction, bit rate reduction and learning-driven rate-distortion optimization. Dimension reduction aims to introduce bottleneck layers at the splitting point of DNNs to compress the dimension of semantics, which can be further classified as reconstruction-based and task-oriented compression. Bit rate reduction leverages either lossless or lossy coding techniques to reduce the data size that needs to be transmitted. The taxonomy of SEC is depicted in Figure 4.
@@ 3.1.1 System Optimization
The system optimization problem such as Eqn. 1 is often NP-hard, hence requiring advanced approaches to find the optimal solution. Based on the optimization approach, literature can be categorized as heuristic search [64], solver-based [51], relaxation-based [49] as well as DRL-based approaches [52]. Based on the partitioning graph, it can also be classified as sequential-based [65] and DAG-based [66] partitioning.
@@ 3.1.2 Semantic Compression
The transmission overhead in SEC can be reduced by compressing the dimension and bit-width of semantic representations. For example, a general approach to add an auto-encoder to encode the original intermediate representation to a lower space on mobile devices and reconstruct the representation at the edge. Also, knowledge distillation [55], NAS [56] and neural network pruning [57] to manually create a bottleneck at the splitting point in order to reduce the dimensionality of latent representations. On the other hand, literature in SEC leverages different source coding approaches such as lossless coding (e.g. run-length coding [68], entropy coding [58]) and lossy coding (e.g. JPEG [69], quantization [6]) to further minimize the the average bit-length of latent representations. Recently inspired by the neural image compression [70], learning-driven rate-distortion optimization approaches are adopted to SEC [58]. The key idea is to optimize DNN based on the rate-distortion metric during training. After training, the learned latent representation can achieve minimum rate with entropy coding under certain distortion level. In addition, dynamic feature compression as an emerging technique has been studied recently to achieve wireless channel adaptive compression of latent representations [71].
@@ 3.2.1 Resource-Aware Partitioning
Sequential Partitioning. Eshratifar et al. [49] investigated layer-wise splitting to minimize the energy consumption of mobile devices and end-to-end latency. It first profiled the energy consumption and execution latency of each DNN layer during run-time. The optimal DNN partitioning is then converted into integer linear programming problem and lagrange relaxation based method [74] is used to constitute a lookup table for a different set of energy and networking constraints. Similarly, Zawish et al. [63] optimized the partitioning under various computing resource constraints (number of FLOPs), energy constraints and bandwidth constraints. Gao et al. [51] benchmarked layer-wise processing time on mobile and edge devices of different hyper parameters such as convolutional kernel size and data size. After benchmarking, a processing time estimation approach based on polynomial regression is developed. A mixed integer linear programming problem is formulated to minimize the end-to-end latency and solved with an opensource solver. Gao et al. [65] further considered the DNN partitioning in a multi-user scenario where multiple mobile devices will connect to the same edge device. The study jointly optimize the partitioning point and task scheduling to minimize inference latency, energy consumption and service fee paid to the server based on the aggregative game theory. Ding et al. [75] considered a privacy-preserving scenario for resource-limited DNN partitioning. The proposed approach first collects per-layer computation and memory cost to build a set of valid splitting points. Then a distance correlation metric [76] is leveraged during training to reduce the information leakage of latent representations. Li and Bi [52] extended the DNN partitioning to the multi-user cooperative inference system. The splitting and resource allocation are jointly optimized by transforming the problem into a mixed integer non-linear programming. A DRL algorithm is used to select splitting point to minimize energy consumption while a convex optimization algorithm is used for resource allocation. Jiang et al. [77] jointly optimized the partitioning and early exit of multi-branch DNNs to minimize the energy consumption. Singhal et al. [78] focused on deploying multi-branch DNNs across mobile-edge-cloud systems. The optimization problem is modeled using a graph model to minimize the layer-to-layer energy cost. Jung et al. [79] used a heuristic search to optimize the energy consumption and over-all latency in a multi-user scenario where multiple mobile devices and a single server performs continuous inference.
@@
DAG Partitioning. While above-mentioned literature considered sequential splitting, Li et al. [66] investigated the DAG-based splitting and inference parallelism to accelerate the multiple inference requests. It involves partitioning the DNN model and offloading parts to cloudlets while using multi-threading to meet inference delay requirements. Duan and Wu [80] categorized DNN graphs as tree-structure and path-independent structure and proposed different algorithms regarding the type of DAG to optimize end-to-end latency under resource constraints. Similarly, Wang et al. [64] partitioned DNN based on its DAG structure to minimize latency. The problem is modeled with mixed integer programming which is proven np-hard and can be solved with heuristic search. Duan and Wu [81] extended the partitioning problem for multiple DNN tasks and proposed to minimize overall latency by jointly optimizing splitting point and task scheduling. Liu et al. [82] leveraged a greedy algorithm that can handle both chain and DAG structure to minimize the end-to-end latency.
@@ 3.2.2 Dynamic Model Partitioning
Sequential Partitioning. Neurosurgeon [62] is one of the earliest work considering dynamic model partitioning in different bandwidth and energy constraints. It first leverages a regression model to estimate the computing load (number of FLOPs) of different DNN layers. Then, a greedy search approach is leveraged to select the splitting point in run-time to minimize the end-to-end latency or energy consumption of mobile devices. Li et al. [83] combined DNN splitting with early exits and optimized the splitting point and exit point jointly in the changing networking environment where the bandwidth varies drastically based on the bayesian online change point detection [84]. Mehta and Shorey [85] proposed to split DNN dynamically based on the changing computation resources to minimize the overall transmission consumption. However, [85] does not consider the dynamic networking condition. Karjee et al. [86] optimize the splitting point to minimize the computation latency in the dynamic computing resource constraints. In order to further optimize the communication latency, a dynamic network switching approach by leveraging both DRL and rule-based selection are proposed to change the wireless techniques such as bluetooth, cellular and Wi-Fi based on the network traffic. Lim and Kim [87] proposed to minimize the end-to-end latency using Lyapunov optimization. By modeling the dynamics in different time steps as a Lyapunov drift, [87] utilized an real-time splitting approach to jointly determine the DNN splitting point and the dynamic voltage and frequency scaling of the GPU on devices. Bakhtiarnia et al. [67] first profiled natural bottlenecks in off-the-shelf DNNs where the output is smaller than the original input data. Then, the DNN is dynamically allocated on mobile and edge device aiming at minimizing the end-to-end latency. Ko et al. [88] investigated the dynamic model partitioning in the distributed severless edge scenario where different head and tail models are running in containers with different computing status. The optimization problem is modeled with a Markov decision process and linear programming is used to choose the optimal container to minimize the inference latency. Nagamatsu and Hara-Azumi [89] utilized mixed-precision NAS [90] to optimize the bit rate of latent representations for different workloads. As the bit rate can be dynamically changed, the DNN is dynamically partitioned based on greedy search in order to minimize the inference time. Lee et al. [91] proposed a two-stage wireless channel adaptive model partitioning addressing the energy and memory constraint of devices. The splitting point is first determined using an off-line exhaustive search without considering the wireless channel. After deployment, the model can be adjusted online to minimize inference latency according to wireless channel conditions. This approach reduced the search space in the online adjustment phase making it easy to deploy in real time. However, the limited adjustment points can result in sub-optimal latency minimization. Liu et al. [92] leveraged multi-agent DRL to optimize the partitioning, task offloading and resource allocation in dynamic vehicular networks. Vehicles in the network act as independent agents to achieve distributed inference in a decentralized manner. Qi et al. [93] leveraged DRL to jointly optimize the model compression and model partitioning in varying network conditions.
@@
DAG Partitioning. Along with chain-based splitting, literature also investigate DNN partitioning based on their DAG structure. For example, Hu et al. [94] investigated the dynamic DAG partitioning in different DNN workload scenarios to minimize the latency in low network traffic scenario and maximize the overall throughput in heavy network traffic scenario. Zhang et al. [95] introduced a dynamic DAG partitioning algorithm that can adjust the splitting point to its nearly vertex in the run time. Liang et al. [50] further extended the work in [94] to a computing resource constrained scenario. Li et al. [96] proposed an adaptive DAG-based splitting to optimize the latency and resource utilization in the multi-user scenario. Nagamatsu et al. [97] applied NAS to optimize potential splitting points and dynamically selected splitting points based on the network condition.
@@
[77] Image classification Sequential minimize energy consumption N/A
[78] Image classification Sequential minimize energy consumption, select early exit N/A
[79] Image classification Sequential optimize latency and energy for multiple devices N/A
[82] Image classification Sequential and DAG minimize latency N/A
@@
[92] Traffic sign detection Sequential minimize latency Yes
[93] Image classification Sequential balance latency and model compression Yes
@@
[97] Image classification DAG minimize latency Yes
3.2.3 Research Challenges
Model Dependencies. One significant challenge in system optimization is finding the accurate computation characteristics–such as latency, energy consumption, and computing memory–of each possible partitioning point. Most existing literature relies on an exhaustive performance analysis for every layer in DNNs [49, 51, 65, 75]. While this approach provides accurate profiling, it makes the optimization model-specific. For different DNN architectures such as CNNs and transformers that have distinct computation and memory usage patterns, even when they have the same computation complexity (number of FLOPs), the latency can be significantly different. This makes optimization algorithms designed for a specific model potentially non-generalizable to others.
Hardware Dependencies. Hardware with different computing capabilities can produce varying latency for identical computation patterns [98]. Only profiling layer-wise execution of DNNs on a single device may fail to generalize to heterogeneous platforms. different platforms often use different DNN compilers (e.g., TensorFlow, PyTorch, ONNX) with varying compatibility requirements. Creating partitioning frameworks that can effectively accelerate DNNs across heterogeneous infrastructures presents significant research challenges [82].
Optimization Complexity. The search space for optimal partitioning is growing exponentially with an increasing DNN model complexity (e.g., DNNs with hundreds of layers), making it computationally intractable to evaluate all possible configurations. Approaches that rely on heuristics or approximation algorithms [64, 79, 89] can reduce the complexity but remain challenging in finding globally optimal solutions. Complex optimization algorithms [52, 99, 65] provide better optimized performance but result in high latency, preventing real-time adaptation in dynamic environments.
Reconfiguration Cost. Network conditions, hardware availability, and workload characteristics change during distributed inference. Static partitioning decisions made at deployment time often become suboptimal, but dynamic repartitioning introduces challenges in maintaining hardware computing status [88]. In addition, most dynamic optimization methods work in a centralized manner [62, 89, 85]. It requires additional orchestration to synchronize the deployment on both mobile and edge devices which result in extra communication and computation overhead.
3.3 Advancements in Semantic Compression
@@ 3.3.1 Dimension Reduction
On the other hand, inspired by DNN compression, the task-oriented bottleneck proposes to directly squeeze the network architecture and train the modified model with task-oriented loss (e.g. cross-entropy loss). Model compression approaches such as knowledge distillation [55], NAS [56] and DNN pruning [57] are often leveraged to achieve superior compression performance while preserving accuracy. For example, Eshratifar et al. [4] modified the size of the early layers in the DNN to decrease computational load on mobile devices. The bottleneck is introduced in the middle of the modified architecture and the model is trained with cross entropy loss. However, it suffers performance loss as introducing bottleneck in early layers can result in loss of task-oriented information. As such, Matsubara and Levorato [108] proposed to train the bottleneck using knowledge distillation. Similarly, Dong et al. [109] optimized the feature compression design using NAS while Guo et al. [110] utilized pruning to reduce the bottleneck size. Mendula et al. [111] use knowledge distillation and FP16 quantization to reduce the communication overhead for object detection tasks.
@@
[118] N/A adaptive bitwidth based on attn map No
[119] auto-encoder adaptive bitwidth and entropy coding No
[120] distillation entropy coding Yes
@@
[126] reconstruction entropy coding Yes